# From Reals to Logic and Back: Inventing Symbolic Vocabularies, Actions and Models for Planning from Raw Data

### v2

@@ 2 Preliminaries
A robot planning domain is defined as a tuple $\langle{\mathcal{O}},{\mathcal{T}},{\mathcal{X}},{\mathcal{A}}\rangle$ where $\mathcal{O}$} is a set of objects, $\mathcal{X}$} is a state space, and $\mathcal{A}$} is an uncountably infinite set of primitive deterministic actions.
@@
A robot planning problem is defined as a tuple $\langle x_{i},{\mathcal{X}}_{g}\rangle$ where $x_{i}\in{\mathcal{X}_{\text{free}}}$ is an initial state and ${\mathcal{X}}_{g}\subseteq{\mathcal{X}_{\text{free}}}$ is a set of goal states.
@@
$\bar{{\mathcal{A}}}$ defines the set of symbolic lifted actions defined using the set of lifted predicates $\mathcal{P}$}. Each abstract action $\bar{a}\in\bar{{\mathcal{A}}}$ is parameterized with typed symbolic parameters. Each action $\bar{a}$ is defined as a tuple $\langle\emph{pre}_{\bar{a}},\emph{eff}_{\bar{a}}\rangle$. Here, $\emph{pre}_{\bar{a}}$ is a conjunctive formula of parameterized predicates from the set of predicates $\mathcal{P}$}. $\emph{eff}_{a}$ is the effect of the action $\bar{a}$. It is defined as a tuple $\emph{eff}_{\bar{a}}=\langle\emph{add}_{\bar{a}},\emph{del}_{\bar{a}}\rangle$ where $\emph{add}_{\bar{a}}$ is a set of predicates that are added to the state and $\emph{del}_{\bar{a}}$ is a set of predicates that are removed from the state when the action $\bar{a}$ is executed. Similarly to the predicates, an action $\bar{a}$ can be grounded using the objects yielding a grounded action $\bar{a}^{\prime}$. This also generates grounded precondition $\emph{pre}_{\bar{a}^{\prime}}$ and grounded effect $\emph{eff}_{\bar{a}^{\prime}}$. A grounded action $\bar{a}^{\prime}$ is applicable in a state $s$ if and only if $\emph{pre}_{\bar{a}^{\prime}}\models s$. Formally, every deterministic grounded action $\bar{a}^{\prime}\in\bar{{\mathcal{A}}}^{\prime}$ defines a function $\bar{a}^{\prime}:s_{i}\mapsto s_{j}$ that maps each symbolic state $s_{i}$ to the resulting state $s_{j}$.
@@ 3 Our approach
Let $T_{\emph{train}}$ be a set of training problems and ${\mathcal{D}}_{\emph{train}}$ be a set of demonstrations that successfully solve the training problems. We define the abstraction learning problem as learning $1)$ a predicate vocabulary $\mathcal{P}$}, $2)$ a set of high-level actions $\bar{{\mathcal{A}}}$, and $3)$ a set of generators $\Gamma$ for each learned predicate.
@@ 1 Use $\xi$ to compute trajectories with relative poses of each object
3 $\mathcal{R}$$\leftarrow$ discover_relations($\Psi$$\mathcal{O}$}, $\mathcal{T}$})
4 $\tilde{{\mathcal{R}}}\leftarrow$ discover_auxilary_relations($\mathcal{R}$});
@@ 6 $\bar{{\mathcal{A}}}\leftarrow$ invent_actions($\mathcal{D}$}, $\mathcal{P}$});
7
[…]
@@ 3.1.2 Represetning Invented Critical Regions as Relations
Relations invented by our approaches can be easily translated into PDDL representation (or any other representation). For a given pair of object types $\tau_{i},\tau_{j}\in{\mathcal{T}}$, let $\Psi_{ij}\subset\Psi$ be the set of critical regions. For each relational critical region $\psi_{k}\in\Psi_{ij}$, each relation $r_{ij}$ can be translated into a binary predicate $\mathtt{(p^{\psi_{k}}_{ij}\;?y_{i}\;?y_{j})}$ where $\mathtt{?y_{i}}$ is a typed parameter of type $\tau_{i}$ and $\mathtt{?y_{j}}$ is a typed parameter of type $\tau_{j}$. Similarly, each auxiliary predicate $\tilde{r}_{ij}$ can be translated to a unary predicate $\mathtt{(\tilde{p}^{\psi_{k}}_{ij}\;?y_{i})}$ where $\mathtt{?y_{i}}$ is a typed parameter of type $\tau_{i}$. Let $\mathcal{P}$} be a set of predicates for all relations (line 1).
@@
1 $\bar{{\mathcal{D}}^{\prime}}_{\emph{train}}\leftarrow$ get_abstract_demonstrations(${\mathcal{D}}_{\emph{train}}$$\mathcal{P}$});
@@ 15  $\emph{pre}\leftarrow\cap_{s\in S_{i}}s$;
16  $\emph{pre}\leftarrowprune$_precondition(pre);
17  $\emph{param}\leftarrow$ extract_params($S_{i}\rightarrow S_{j}$);
18  $\bar{{\mathcal{A}}}$.add(create_action($\emph{param},\emph{pre},\emph{eff}$));
@@ 3.2.2 Learning Symbolic Action Model
1
[…]
3 $G\leftarrow$ create_relation_graph($x_{g}$$\mathcal{P}$});
@@ 8  $s_{g}\leftarrow$ create_symbolic_goal($G$);
9  $\bar{\Pi}\leftarrow$ compute_kk_symbolic_plans($\mathcal{M}$}, $\mathcal{O}$}, $s_{i}$, $s_{g}$, $k$);
@@ 4 Empirical Evaluation
We present the salient aspects of our implementation, setup, and observations here. Our empirical evaluation is designed to answer the following key questions: $1)$ Are the learned abstractions sound and generalizable to unseen complex planning problems?; $2)$ Are the learned abstractions transferrable to different similar robots?; and $3)$ How close the learned abstractions are to human intuition?
@@ 4.1 Analysis of Results
Table 1: Detailed statistics about the empirical evaluation and invented abstractions. The success rate is an average of $50$ independent test tasks with 55 random seeds.
@@
Table 2: Ablation study of our approach with decreased training demonstrations and comparison with baseline approaches. Success rate for our approach and baselines averaged over $10$ different unseen test tasks and 55 random executions. The percentages represent the fraction of training demonstrations used for learning the initial state and action abstractions.
@@
Table 2 also illustrates the number of training demonstrations utilized for different evaluations of our approach, emphasizing the scalability and generalizability of our method even with limited data. With a modest number of training demonstrations, our approach outperformed the baselines in complex problems. It successfully tackled most tasks with abstractions learned in a few-shot manner. Specifically, our approach achieved a 100100% success rate in tasks involving building structures with Keva planks and packing cans in a box, needing only $50$ trajectories. This underscores the data-efficiency of our approach and its ability to generalize effectively from a small number of demonstrations. Despite the non-trivial nature of learning abstractions for a mobile manipulation problem, our method efficiently solved $100\%$ of these tasks in a cafe setting. However, for more intricate trajectories, such as grasping cans from different sides while positioning the robot on various sides of the table, it required a relatively higher number of trajectories ($500$). These training demonstrations, as noted above, also include $\sim 50\%$ randomly demonstrations that do not achieve the goal.
@@ Appendix B Example of Learning Actions
  • $S_{1}:\{$(table-can1 ?table ?can), (can-gripper1 ?can ?gripper), (base-gripper0 ?base ?gripper), (base-table1 ?base ?table)}\}.
  • $S_{2}:\{$(table-can1 ?table ?can), (can-gripper2 ?can ?gripper), (base-gripper0 ?base ?gripper), (base-table1 ?base ?table)}\}.
  • $S_{3}:\{$(table-can0 ?table ?can), (can-gripper2 ?can ?gripper), (base-gripper0 ?base ?gripper), (base-table1 ?base ?table)}\}.
  • $S_{4}:\{$(table-can0 ?table ?can), (can-gripper2 ?can ?gripper), (base-gripper1 ?base ?gripper), (base-table1 ?base ?table)}\}.
@@
Learning preconditions involve taking intersection of states in which all the actions in the same cluster were executed. Here $S_{1}$ to $S_{3}$ mentioned below will remain the same for the three clusters. For e.g., precondition of $C_{12}=\{$(table-can1 ?table ?can), (can-gripper1 ?can ?gripper), (base-gripper0 ?base ?gripper), (base-table1 ?base ?table)}\}. Alg. 2 will prune out (base-table1 ?base ?table) from the precondition as (i) it is unchanged between $S_{1}$ and $S_{2}$, and (ii) none of its parameters (?base and ?table) are part of any other predicate that is changed. Using this, the precondition for each action will be:
  • $\emph{pre}(C_{12})=\{$(table-can1 ?table ?can), (can-gripper1 ?can ?gripper), (base-gripper0 ?base ?gripper)}\}.
  • $\emph{pre}(C_{23})=\{$(table-can1 ?table ?can), (can-gripper2 ?can ?gripper), (base-gripper0 ?base ?gripper), (base-table1 ?base ?table)}\}.
  • $\emph{pre}(C_{34})=\{$(table-can0 ?table ?can), (can-gripper2 ?can ?gripper), (base-gripper0 ?base ?gripper), (base-table1 ?base ?table)}\}.
@@ Appendix C Test Environments
[Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image][Uncaptioned image]

### v3

# From Reals to Logic and Back: Inventing Symbolic Vocabularies, Actions, and Models for Planning from Raw Data

@@ 2 Preliminaries
A robot planning domain is defined as a tuple $\langle{\mathcal{O}},{\mathcal{T}},{\mathcal{X}},{\mathcal{A}}\rangle$ where ${\mathcal{O}}$ is a set of objects, ${\mathcal{T}}$ is a set of object types, ${\mathcal{X}}$ is a state space, and ${\mathcal{A}}$ is an uncountably infinite set of primitive deterministic actions.
@@
[…]
@@
[…]
@@ 3 Our approach
Let $T_{\emph{train}}$ be a set of training problems and ${\mathcal{D}}_{\emph{train}}$ be a set of demonstrations that successfully solve the training problems. We define the abstraction learning problem as learning 1)1)1 ) a predicate vocabulary ${\mathcal{P}}$, 2)2)2 ) a set of high-level actions $\bar{{\mathcal{A}}}$, and 3)3)3 ) a set of generators $\Gamma$ for each learned predicate.
@@ 1 Use $\xi$ to compute trajectories with relative poses of each object
3 ${\mathcal{R}}\leftarrow$ discover_relations($\Psi$,${\mathcal{O}}$, ${\mathcal{T}}$)
4 $\tilde{{\mathcal{R}}}\leftarrow$ ${\mathcal{R}}$);
@@ 6 $\bar{{\mathcal{A}}}\leftarrow$ invent_actions(${\mathcal{D}}$, ${\mathcal{P}}$);
7 ${\mathcal{M}}\leftarrow$ generate_PDDL(${\mathcal{T}}$,${\mathcal{O}}$${\mathcal{R}}$,$\bar{{\mathcal{A}}}$);
@@ 3.1.2 Represetning Invented Critical Regions as Relations
[…]
@@
1 $\bar{{\mathcal{D}}^{\prime}}_{\emph{train}}\leftarrow$ get_abstract_demonstrations(${\mathcal{D}}_{\emph{train}}$,${\mathcal{P}}$);
@@ 15       $\emph{pre}\leftarrow\cap_{s\in S_{i}}s$;
[…]
@@ 3.2.2 Learning Symbolic Action Model
1 $s_{i}\leftarrow$ get_abstract_state($x_{i}$);
2 $G\leftarrow$ create_relation_graph($x_{g}$,${\mathcal{P}}$);
@@ 7       $s_{g}\leftarrow$ create_symbolic_goal($G$);
8       $\bar{\Pi}\leftarrow$ compute_$k$${\mathcal{M}}$, ${\mathcal{O}}$, $s_{i}$, $s_{g}$, $k$);
@@ 4 Empirical Evaluation
We present the salient aspects of our implementation, setup, and observations here. Our empirical evaluation is designed to answer the following key questions: 1)1)1 ) Are the learned abstractions sound and generalizable to unseen complex planning problems?; 2)2)2 ) Are the learned abstractions transferrable to different similar robots?; and 3)3)3 ) How close the learned abstractions are to human intuition?
@@ 4.1 Analysis of Results
Table 1: Detailed statistics about the empirical evaluation and invented abstractions. The success rate is an average of $50$ independent test tasks with $5$ random seeds.
@@
Table 2: Ablation study of our approach with decreased training demonstrations and comparison with baseline approaches. Success rate for our approach and baselines averaged over $10$ different unseen test tasks and $5$ random executions. The percentages represent the fraction of training demonstrations used for learning the initial state and action abstractions.
@@
Table 2 also illustrates the number of training demonstrations utilized for different evaluations of our approach, emphasizing the scalability and generalizability of our method even with limited data. With a modest number of training demonstrations, our approach outperformed the baselines in complex problems. It successfully tackled most tasks with abstractions learned in a few-shot manner. Specifically, our approach achieved a $100$% success rate in tasks involving building structures with Keva planks and packing cans in a box, needing only $50$ trajectories. This underscores the data-efficiency of our approach and its ability to generalize effectively from a small number of demonstrations. Despite the non-trivial nature of learning abstractions for a mobile manipulation problem, our method efficiently solved $100\%$ of these tasks in a cafe setting. However, for more intricate trajectories, such as grasping cans from different sides while positioning the robot on various sides of the table, it required a relatively higher number of trajectories ($500$). These training demonstrations, as noted above, also include $\sim 50\%$ randomly demonstrations that do not achieve the goal.
@@ Appendix B Example of Learning Actions
[…]
@@
[…]